OpenStudy (anonymous):

Find the sum. 1+4+16...+65,536

OpenStudy (anonymous):

its a gp series

OpenStudy (anonymous):

since common ratio is same

OpenStudy (anonymous):

whats the sum?

OpenStudy (anonymous):

\[4^0 + 4^1+4^2 + \cdots + 4^8 \]

OpenStudy (kropot72):

The r th term is:\[ax ^{r-1}\] where a is the first term and x is the common ratio. \[65536=4^{r-1}\] Solving gives r = 9 So 65536 is the 9th term of the geometric progression the sum of n terms is:\[\frac{a(x ^{n}-1)}{x-1}=\frac{4^{9}-1}{4-1}=87381\]

OpenStudy (anonymous):

thank you!

OpenStudy (kropot72):

You're welcome :)